The manager of a publishing company plans to give a $26,000 bonus to the top 11 percent, $14,000 to the next 29 percent, and $6,000 to the next 10 percent of sales representatives. If the publishing company has a total of 180 sales representatives, what is the expected bonus that the company will pay?

Expected Bonus:_______
Answer #1

Expected bonus for single sales representative = 26000 * 0.11 + 14000 * 0.29 + 6000 * 0.1 = $ 7520

Expected bonus the company will pay to 180 sales representative = 180 * 7520 = $1353600 (ans)